the graph of y=-3x+10 and y=-2/3x+3 meet at the point where
Answer by mananth(16946)

3 x + 1 y = 10 .............1
0.67 x + 1 y = 3 .............2

multiply (1)by -1
Multiply (2) by 1
-3 x -1 y = -10
0.67 x + 1 y = 3
Add the two equations
-2.33 x = -7
/ -2.33
x = 3
plug value of x in (1)
3 x + 1 y = 10
9 + 1 y = 10
1 y = 10 -9
1 y = 1
y = 1
(3,1)